OpenStudy (mirandaregina):

How do I evaluate 8sin(30) - 6cos(60) ?

OpenStudy (trojanpoem):

Cos(60) = sin(30) = 0.5 8 * 0.5 - 6 * 0.5 = 0.5(8-6) = 0.5 * 2 = 1

OpenStudy (mirandaregina):

Thank you @TrojanPoem !

OpenStudy (trojanpoem):

You are welcome.